OpenStudy (anonymous):

ok guys i am having problems with rational numbers and it says simplify 7x+9/7x+3 I am thinking the answer is 7x+3/7x+1 is this correct?

OpenStudy (mathteacher1729):

\[\huge \frac{7x+9}{7x+3}\] is simplified. There's nothing more to do here.

OpenStudy (anonymous):

No, that answer is not right. You can simplify at best as: (7x+3+6)/(7x+3) = 1 + 6/(7x+3). I not sure if this is more simplified than original.

OpenStudy (anonymous):

0 great

OpenStudy (anonymous):

i guess i just dont get it

OpenStudy (anonymous):

(7x+9)/(7x+3) can't be simplified like the way you did. You can't simply divide 9 by 3. You have to treat (7x+9) as a single expression and similarly (7x+3).

OpenStudy (mathteacher1729):

GT -- you make a good point. I should have been clearer in saying that you cannot just "cancel the 7s" or "cancel the 9/3 to 3/1") because those are not like terms in this rational expression. Thank you for the clarification.

OpenStudy (anonymous):

If the problem was (21x+9)/(7x+3), then you can factor 3 from the numerator and get 3*(7x+3)/(7x+3) = 3.